# LaT-PFN: A Joint Embedding Predictive Architecture for In-context Time-series Forecasting

### v1

@@ Abstract
We introduce LatentTimePFN (LaT-PFN), a foundational Time Series model with a strong embedding space that enables zero-shot forecasting. To achieve this, we perform in-context learning in latent space utilizing a novel integration of the Prior-data Fitted Networks (PFN) and Joint Embedding Predictive Architecture (JEPA) frameworks. We leverage the JEPA framework to create a prediction-optimized latent representation of the underlying stochastic process that generates time series and combines it with contextual learning, using a PFN. Furthermore, we improve on preceding works by utilizing related time series as a context and introducing an abstract time axis. This drastically reduces training time and increases the versatility of the model by allowing any time granularity and forecast horizon. We show that this results in superior zero-shot predictions compared to established baselines. We also demonstrate our latent space produces informative embeddings of both individual time steps and fixed-length summaries of entire series. Finally, we observe the emergence of multi-step patch embeddings without explicit training, suggesting the model actively learns discrete tokens that encode local structures in the data, analogous to vision transformers.
@@ 1 Introduction
In this paper, we introduce LatentTimePFN (LaT-PFN), a foundational Time Series model designed to address zero-shot forecasting by combining the Prior-data Fitted Networks (PFN) and Joint Embedding Predictive Architecture (JEPA) frameworks. LaT-PFN is trained exclusively on a novel synthetic data generation method, allowing us to encode expert knowledge directly in the training data. Furthermore, LaT-PFN learns to perform in-context learning in latent space by approximating the Posterior Predictive Distribution (PPD). We implement this meta-learning approach by enabling the model to learn zero-shot forecasting at test time, with user-provided time series functioning as exemplary context. We train LaT-PFN using an abstract time dimension, which consolidates a wider range of time-frequency patterns, facilitating meta-learning. Additionally, this improves the model’s versatility and accuracy, whilst significantly reducing computational requirements. Following the JEPA methodology, LaT-PFN separates predicting and decoding, which are optimized independently. As such, LaT-PFN is trained to predict the next latent state and exploits a system identification loss as a regularization term, to independently improve the quality of its embedding space.
@@
  • We introduce a novel architecture, LaT-PFN, which introduces a combination of the PFN and JEPA frameworks for zero-shot predictions of univariate time series in latent space.
  • We introduce a novel synthetic prior for simulating context-aware time series data. Specifically, our prior allows for the creation of abstract-time synthetic contexts, defined as collections of time series, similar enough to enable LaT-PFN to approximate the PPD.
@@ 2.1 PFN: Prior-data Fitted Networks
The PPD is typically defined as the probability distribution $P(y_{*}|x_{*},\mathcal{D})$, with $\mathcal{D}=\{(x_{i},y_{i})\}^{N}$ being a context dataset. The PFN framework works by approximating this distribution:
@@
Here $Q$ is a transformer [54] parametrized by $\theta$, $S(.)$ is a simulation engine, and $Pr(.)$ is a prior distribution over instances of the simulation parameters $\psi\in\Psi$. Examples of the relation between features and targets are provided by the tuples $(x_{i},y_{i})$ within $\mathcal{D}$, which are then passed through a transformer encoder [54]. The held-out features $x_{*}$ are passed through the decoder, with a diagonal-only (independent) target mask, allowing only attention over the encoder context (Figure 1). Finally, a head and a cross-entropy loss are applied. The network $Q$ is trained to perform meta-learning on any dataset $\mathcal{D}_{j}$ [41]. The authors prove that this is a valid approximation to the PPD – see Appendix B. Since its introduction, others have applied the PFN framework to various tasks [29, 17, 4, 23]. Notably, TabPFN [29] solves small tabular datasets, zero-shot. Additionally, ForecastPFN [17] introduced this methodology to time-series forecasting thanks to a novel synthetic prior.
@@ 2.2 JEPA: Joint Embedding Predictive Architecture
The Joint Embedding Predictive Architecture (JEPA) [35] framework aims to create strong embeddings of any modality, to efficiently handle complex relationships within data. The key concept is to predict state transitions $S_{t}\rightarrow S_{t+1}$ fully in latent space, in order to create embeddings that are not burdened by unnecessary detail. Figure 2 illustrates the core components, which define a sequence of transformations aimed at processing inputs within latent space. Initially, an input embedder $\bar{x}=E_{x}(x)$ converts the input into latent space.
@@
This is followed by a predictor $\hat{\bar{y}}=P(\bar{x},z)$ which estimates the next latent state, using the transformed input and a variable $z$, intended to capture relationships not observable in the data space. A target embedder $\bar{y}=E_{y}(x,y)$ then maps the target, taking into account both $x$, $y$ and their the relationship. A criterion, typically a Mean Squared Error (MSE) loss $C(\hat{\bar{y}},\bar{y})$, evaluates the latent prediction. Finally, a generative decoder, $\hat{y}=D(\hat{\bar{y}})$, decodes the latent representations back to the data space.
The JEPA framework has since been applied to image and video data [5, 6]. The target encoder $y_{L}=E_{y}(x,y)$ is updated as an exponential moving average (EMA) of the input encoder $E_{x}$ and followed by a stop gradient, preventing collapse by learning an identity mapping [35, 5]. Other JEPA works include Motion [7], Audio [22], Point-cloud [46] and EEG transfer [27].
@@ 3.1 Problem Statement
Figure 3: Problem statement of a time series forecasting PFN
@@
Where $\mathcal{D}$ is a context dataset with $N$ examples, for $\mathcal{D}\in\mathbb{R}^{N\times S\times F}$, sequence dimension $S$, example dimension $N$ and features $F$. This differs from the ForecastPFN formulation, which defines the context $\mathcal{D}\in\mathbb{R}^{S\times F}$ as simply the history of the target time series [17]. This interpretation, although elegant, does not allow a context for the TSF task definition and is simply a regular forecasting task with simulated data. See Figure 3 for a visualization.
To complement this notation, let $\bar{x}=\bar{v}_{0:t}$ and latent target $\bar{y}=\bar{v}_{t+1:H}$ be the embedded versions of history $x=v_{0:t}$ and target $y=v_{t+1:H}$ respectively. Similarly, let $\hat{\bar{y}}=\hat{\bar{v}}_{t+1:H}$ be the latent forecast, and $\hat{y}=\hat{v}_{t+1:H}$ be the forecast in data space.
3.2 Abstract Time Dimension
Unlike [17], instead of using absolute time units (year, month, etc.), we map time to a fixed abstract interval $T_{0}:T_{H}$ relative to the forecast start $t$ of each series, with a variable sampling rate – while ensuring the absence of any data leakage. This allows us to map context time series and past segments of the held-out series onto the same interval, independent of their actual place in history. The approach aligns well with the idea of using simulated data to learn Bayesian inference, as it creates a universal and consistent time dimension in which the model can more readily learn general temporal patterns. Furthermore, reducing the number of temporal degrees of freedom that require training, leads to a more efficient convergence and significantly decreases the computational demands for training.
@@ 3.3 An Architecture for Latent In-context Forecasting
Figure 4: The LaT-PFN Architecture. The context is embedded in fixed-length series-vectors. These are fed into the PFN Predictor transformer, with the embedded held-out history prompts, using cross-attention. The latent predictions are compared to the latent target, then decoded with a stopgradient, and compared to real targets. Finally, we apply a supervised regularisation on simulation parameters.
@@
The embedder $E(.)$ is an eight-layer dilated Mobilenet1D [30]. We empirically found convolutions to be superior to attention over the sequence dimension, in line with [63, 38]. The component is responsible for feature extraction along the sequence dimension. First, we create the context embeddings $\bar{\mathcal{D}}$ and the latent target $\bar{y}$ with time $T$ and value $V$ features. For the latent target specifically, we use an EMA of the embedder parameters, apply a stop gradient, and select only the future timesteps:
@@
The predictor $PFN_{\theta}(.)$ forecasts the next latent state, given the prompt and context. We apply learned average pooling $AVG_{\theta}$ over the sequence dimension of the embedded context $\bar{\mathcal{D}}$, by cross-attention with a learned vector $\vec{q}\in\theta$. These are then provided to the encoder of the latent PFN predictor 4, with self-attention over the examples dimension.
$\bar{\eta}\hskip 5.0pt=\hskip 5.0ptAVG_{\theta}(\bar{\mathcal{D}})$
In order to apply 2D attention over the held-out series, we flatten the examples dimension $N$ and sequence dimension $S$ of the prompt $\bar{z}$. Next, we pass it to the decoder, whilst ensuring adjacent sequences from held-out series are not dependent on sharing information.
$M\hskip 3.30002pt=\hskip 3.30002ptdiag(1)^{N\cdot S\times N\cdot S},\quad\hat{\bar{y}}\hskip 3.30002pt=\hskip 3.30002ptPFN_{\theta}(\bar{z},\bar{\eta},M),\quad\mathcal{L}_{\text{latent}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{\bar{y}}-\bar{y})^{2}$
@@
The decoder $D_{\theta}(.)$ is a three-layer feedforward network, trained with independent gradients in alignment with the JEPA methodology [35, 5, 6]. We empirically found this to be critical for generating stable embeddings. The decoder utilizes a cross-entropy loss (see Appendix B) to define a categorical distribution over 100 segmented bins of the output space, incorporating label smoothing of 0.01 [50]. This is made feasible by casting the values $V$ to an abstract space through z-normalizing with 2-std. By working in a value-abstract space, the decoder yields an output without making any assumptions on the target distribution – adhering closely to the PFN formulation [41].
$\hat{y}\hskip 3.30002pt=\hskip 3.30002ptD_{\theta}(StopGrad(\hat{\bar{y}})),\quad\mathcal{L}_{\text{decoder}}\hskip 3.30002pt=\hskip 3.30002pt\mathbb{E}_{p(y)}\log p(\hat{y}\>|\>\>\hat{\bar{y}})$
@@
Inspired by Sim2Real research (e.g. [62]), we add a regularization term, to encourage consistency between history and forecast. This component processes the summarised vectors of both context and predicted held-out series, to identify the simulation parameters that generated the context $\mathcal{D}$. This approach further supervises the latent space, directing it to the underlying characteristics of the simulation, such as trend and seasonality. This, in turn, leads to improved forecasting capabilities and stabilizes training [35]. The latent embeddings and predictions are concatenated and averaged, then passed to a head $H_{\theta}(.)$ alongside the embedded context. $H_{\theta}(.)$ learns a multi-target regression towards unit-normalized simulation parameters.
$\hat{y}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002ptH_{\theta}(AVG_{\theta}(\hat{\bar{y}},\bar{x}),\bar{C}),\quad\mathcal{L}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{y}_{\text{si}}-\psi_{j})^{2}$
@@ 3.4 Context-aware Synthetic Prior with Triple Sampling
Figure 6: Decoded fits plotted on the target, with history and target in blue and red scatter respectively, as well as the PPD fore- and backcast mean as blue line with variance, and finally the input example context in green in the background. The fits demonstrate the predictions in latent space have picked up concepts such as seasonality and trend, making it easier to extrapolate after the forecast moment.
Following the PFN example [41, 17, 29], Lat-PFN was trained exclusively on synthetic data. For this work, we adopted and adjusted ForecastPFN’s time series synthetic prior [17]. Like ForecastPFN, we set hyperprior parameters that govern sampling simulation parameters. Unlike ForecastPFN however, our context is defined not by the history of a single time series, but by a collection of example time series. Hence, for most simulation parameters, we introduce a context dimension by employing a three-step sampling strategy:
@@
    Context Parameters $(\alpha_{c})$: Ranges sampled uniformly from the hyperprior to define the limits within an individual context.
@@
    Sub-context Parameters $(\mu)$: Cluster centers as $(\mu_{1},\mu_{2})$ sampled uniformly from the context range, are used to introduce contrast within a context and guide the model to recognize when a context example may be non-informative.
@@
    Series Parameters $(\psi)$: Point estimates normally sampled using a fixed hyperprior variance $\sigma^{*}$, used as simulation parameters in $S(v_{i,0:T}|\psi_{i})$ to generate individual training data series.
This approach ensures balance in inter- and intra-context variance, promoting coherence without trivializing the series and enhancing the model’s ability to generalize in a zero-shot setting. A full description of the synthetic prior and which parameters are used can be found in appendix C.1.

### v2

@@ Abstract
We introduce LatentTimePFN (LaT-PFN), a foundational Time Series model with a strong embedding space that enables zero-shot forecasting. To achieve this, we perform in-context learning in latent space utilizing a novel integration of the Prior-data Fitted Networks (PFN) and Joint Embedding Predictive Architecture (JEPA) frameworks. We leverage the JEPA framework to create a prediction-optimized latent representation of the underlying stochastic process that generates time series and combines it with contextual learning, using a PFN. Furthermore, we improve on preceding works by utilizing related time series as a context and introducing a normalized abstract time axis. This reduces training time and increases the versatility of the model by allowing any time granularity and forecast horizon. We show that this results in superior zero-shot predictions compared to established baselines. We also demonstrate our latent space produces informative embeddings of both individual time steps and fixed-length summaries of entire series. Finally, we observe the emergence of multi-step patch embeddings without explicit training, suggesting the model actively learns discrete tokens that encode local structures in the data, analogous to vision transformers.
@@ 1 Introduction
In this paper, we introduce LatentTimePFN (LaT-PFN), a foundational Time Series model designed to address zero-shot forecasting by combining the Prior-data Fitted Networks (PFN) and Joint Embedding Predictive Architecture (JEPA) frameworks. LaT-PFN is trained exclusively on a novel synthetic data generation method, allowing us to encode expert knowledge directly in the training data. Furthermore, LaT-PFN learns to perform in-context learning in latent space by approximating the Posterior Predictive Distribution (PPD). We implement this meta-learning approach by enabling the model to learn zero-shot forecasting at test time, with user-provided time series functioning as exemplary context. We train LaT-PFN using a normalized time-axis, which consolidates a wider range of time-frequency patterns, facilitating meta-learning. Additionally, this improves the model’s versatility and accuracy, whilst reducing computational requirements. Following the JEPA methodology, LaT-PFN separates predicting and decoding, which are optimized independently. As such, LaT-PFN is trained to predict the next latent state and exploits a system identification loss as a regularization term, to independently improve the quality of its embedding space.
@@
  • We introduce a novel architecture, LaT-PFN, which implements a combination of the PFN and JEPA frameworks for zero-shot predictions of univariate time series in latent space.
  • We introduce a novel synthetic prior for simulating context-aware time series data. Specifically, our prior allows for the creation of synthetic contexts, defined as collections of time series, similar enough to enable LaT-PFN to approximate the PPD.
@@ 2.1 PFN: Prior-data Fitted Networks
The PPD is typically defined as the probability distribution $P(y_{*}|x_{*},\mathcal{D})$, with $\mathcal{D}=\{(x_{i},y_{i})\}^{N}$ being a context dataset. PFNs work by approximating this distribution with network $Q_{\theta}(.)$:
@@
Here $Q$ is a transformer [54] parametrized by $\theta$, $S(.)$ is a simulation engine, and $Pr(.)$ is a prior distribution over instances of the simulation parameters $\psi\in\Psi$. Examples of the relation between features and targets are provided by the tuples $(x_{i},y_{i})$ within $\mathcal{D}$, which are then passed through a transformer encoder [54]. The held-out features $x_{*}$ are passed through the decoder, with a diagonal-only (independent) target mask, allowing only attention over the encoder context (Figure 1). Finally, a head and a cross-entropy loss are applied. The network $Q$ is trained to perform meta-learning on any dataset $\mathcal{D}_{j}$ [41]. The authors prove that this is a valid approximation of the PPD – see Appendix C. Since its introduction, others have applied the PFN framework to various tasks [29, 16, 3, 22]. Notably, TabPFN [29] solves small tabular datasets, zero-shot. Additionally, ForecastPFN [16] introduced this methodology to time-series forecasting thanks to a novel synthetic prior.
@@ 2.2 JEPA: Joint Embedding Predictive Architecture
The Joint Embedding Predictive Architecture (JEPA) [35] framework aims to create a strong representation space to model complex relationships within data, without relying on reconstruction losses or heuristics. The key concept is to predict self-supervised state transitions $S_{t}\rightarrow S_{t+1}$ fully in latent space, yielding embeddings that are not burdened by unnecessary detail. Figure 2 illustrates the core components, which define a sequence of transformations aimed at processing inputs within latent space. Initially, an input embedding $\bar{x}=E_{x}(x)$ is produced by the input embedder.
@@
This is followed by a prediction $\hat{\bar{y}}=P(\bar{x},z)$ constituting the next estimated latent state, using the transformed input $\bar{x}$ and a variable $z$, intended to capture relationships not observable in data space – for example camera angles for computer vision. A target embedding $\bar{y}=E_{y}(x,y)$ is formed, taking into account both $x$, $y$ and their relationship. A criterion, typically a Mean Squared Error (MSE) loss $C(\hat{\bar{y}},\bar{y})$, evaluates the latent prediction. Finally, a generative decoder trained in isolation decodes the latent representations back to the data space $\hat{y}=D(\hat{\bar{y}})$.
The JEPA framework has since been applied to image and video data [4, 5]. The target encoder $y_{L}=E_{y}(x,y)$ is updated as an exponential moving average (EMA) of the input encoder $E_{x}$ and followed by a stop-gradient, preventing collapse by learning an identity mapping [35, 4]. Other JEPA works include Motion [6], Audio [21], Point-cloud [46] and EEG transfer [27].
@@ 3.1 Problem Statement
Figure 3: Our time series forecasting PFN problem statement, compared to preceding works [16, 29]
@@
Where $\mathcal{D}$ is a context dataset with $N$ examples, for $\mathcal{D}\in\mathbb{R}^{N\times S\times F}$, sequence dimension $S$, example dimension $N$ and features $F$. This differs from the ForecastPFN formulation, which defines the context $\mathcal{D}\in\mathbb{R}^{S\times F}$ as the history of the target time series [16]. This interpretation, although elegant, does not allow for a context for the TSF task definition and is essentially a regular TSF approach trained with simulated data. See Figure 3 for a visualization.
To complement this notation, let embedded history $\bar{x}=\bar{v}_{0:t}$ and latent target $\bar{y}=\bar{v}_{t+1:H}$ be the embedded versions of history $x=v_{0:t}$ and target $y=v_{t+1:H}$ respectively. Similarly, let $\hat{\bar{y}}=\hat{\bar{v}}_{t+1:H}$ be the latent forecast, and $\hat{y}=\hat{v}_{t+1:H}$ be the forecast in data space.
3.2 Normalized Abstract Time-axis
Unlike [16], instead of using absolute time units (year, month, etc.), we map time to a fixed normalized interval $T_{0}:T_{H}$ relative to the forecast start $t$ of each series, with a variable sampling rate – while ensuring the absence of any data leakage. This allows us to map context time series and past segments of the held-out series onto the same interval, independent of their actual place in history. The approach aligns well with the idea of using simulated data to learn Bayesian inference, as it creates a universal and consistent time dimension in which the model can more readily learn general temporal patterns. Furthermore, reducing the number of temporal degrees of freedom that require training, leads to a more efficient convergence and decreases the computational demands for training.
@@ 3.3 An Architecture for Latent In-context Forecasting
Figure 4: The LaT-PFN Architecture. The context is embedded in fixed-length series-vectors. These are fed into the PFN Predictor transformer, with the embedded held-out history prompts, using cross-attention. The latent predictions are compared to the latent target, then decoded with a stop-gradient, and compared to real targets. Finally, we apply a supervised regularization on simulation parameters.
@@
The embedder $E_{\theta}$ is an eight-layer dilated Mobilenet1D [30]. We empirically found convolutions to be superior to attention over the sequence dimension, in line with [63, 38]. The component is responsible for feature extraction along the sequence dimension. First, we create the context embeddings $\bar{\mathcal{D}}$ and the latent target $\bar{y}$ with time $T$ and value $V$ features. For the latent target specifically, we use an EMA of the embedder parameters, apply a stop-gradient, and select only the future timesteps:
@@
The predictor $\textit{PFN}_{\theta}$ forecasts the next latent state, given the prompt and context. We apply learned average pooling $\textit{AVG}_{\theta}$ over the sequence dimension of the embedded context $\bar{\mathcal{D}}$, by cross-attention with a learned vector $\vec{q}\in\theta$. These are then provided to the encoder of the latent PFN predictor in Figure 4, with self-attention over the examples dimension.
$\bar{\eta}\hskip 5.0pt=\hskip 5.0pt\textit{AVG}_{\theta}(\bar{\mathcal{D}})$
In order to apply 2D attention over the held-out series, we flatten the examples dimension $N$ and sequence dimension $S$ of the prompt $\bar{z}$. Next, we pass it to the decoder, whilst ensuring adjacent sequences from held-out series are not dependent on sharing information amongst each other.
$M\hskip 3.30002pt=\hskip 3.30002ptdiag(1)^{N\cdot S\times N\cdot S},\quad\hat{\bar{y}}\hskip 3.30002pt=\hskip 3.30002pt\textit{PFN}_{\theta}(\bar{z},\bar{\eta},M),\quad\mathcal{L}_{\text{latent}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{\bar{y}}-\bar{y})^{2}$
@@
The decoder $D_{\theta}$ is a three-layer feedforward network tasked with decoding latent predictions, approximating the PPD in data space (see Appendix C). It is trained with independent gradients in alignment with the JEPA methodology [35, 4, 5]. We empirically found this to be critical for generating stable embeddings upstream. The decoder utilizes a cross-entropy loss to define a categorical distribution over 100 segmented bins of the output space (see Appendix C), incorporating label smoothing of 0.01 [50]. This is made feasible by z-normalizing the values $V$ with 2-std, towards a fixed interval. Consequently, the decoder yields an output without making any assumptions on the target distribution – adhering closely to the PFN formulation [41].
[…]
@@
Inspired by Sim2Real research (e.g. [62]), we add a regularization term, to encourage consistency between history and forecast. This component processes the summarised vectors of both context and predicted held-out series, to identify the simulation parameters that generated the context $\mathcal{D}$. This approach further supervises the latent space, directing it to the underlying characteristics of the simulation, such as trend and seasonality. This, in turn, leads to improved forecasting capabilities and should stabilize training [35]. The latent embeddings and predictions are concatenated and averaged, then passed to a head $H_{\theta}$ alongside the embedded context. $H_{\theta}$ learns a multi-target regression towards unit-normalized simulation parameters.
$\hat{y}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002ptH_{\theta}\big{(}\>\textit{AVG}_{\theta}(\hat{\bar{y}},\bar{x}),\bar{C}\>\big{)},\quad\mathcal{L}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{y}_{\text{si}}-\psi_{j})^{2}$
@@ 3.4 Context-aware Synthetic Prior with Triple Sampling
Figure 6: Decoded fits plotted on the target, with history and target in blue and red scatter respectively, as well as the PPD fore- and backcast mean as a blue line with variance, and finally the input example context in green in the background. The fits demonstrate the predictions in latent space have picked up concepts such as seasonality and trend, making it easier to extrapolate after the forecast moment.
Following the PFN example [41, 16, 29], Lat-PFN was trained exclusively on synthetic data. For this work, we adopted and adjusted ForecastPFN’s time series synthetic prior [16]. Like ForecastPFN, we set hyperprior parameters that govern the sampling of simulation parameters. Unlike ForecastPFN however, our context is defined not by the history of a single time series, but by a collection of example time series. Hence, for most simulation parameters, we introduce a context dimension by employing a three-step sampling strategy:
@@
    Context Parameters $(\alpha_{c})$: Ranges sampled uniformly from the hyperprior to define the limits of an individual context.
@@
    Sub-context Parameters $(\mu)$: Cluster centers $(\mu_{1},\mu_{2})$, sampled uniformly from the context range, are used to introduce contrast within a context and guide the model to recognize when a context example may be non-informative.
@@
    Series Parameters $(\psi)$: Point estimates normally sampled with fixed hyperprior variance $\sigma^{*}$, used as simulation parameters in $S(v_{i,0:T}|\psi_{i})$ to generate individual training series.
This approach ensures balance in inter- and intra-context variance, promoting coherence without trivializing the series and enhancing the model’s ability to generalize in a zero-shot setting. A full description of the synthetic prior and which parameters are used can be found in Appendix D.1.